A multinational company placed orders for printing papers and writing pads. In the first order 5 carton of printing papers and 10 writing pads were ordered and total cost was Rs.3000 and in the second order 10 carton of printing papers and 5 writing pads were ordered and total cost was Rs. 5250. What was the cost of one carton of printing paper
and one writing pad?

Answer by CubeyThePenguin(3113)

P = cost of carton of printing paper
W = cost of writing pad

5P + 10W = 3000
10P + 5W = 5250

Add both equations to get:

15P + 15W = 8250
P + W = 550